OpenStudy (anonymous):

The square root of what = 11? The square root of 225=?

OpenStudy (rulnick):

121 15

OpenStudy (anonymous):

11*11=121 radical 225=15

OpenStudy (anonymous):

How do u know? How do u solve this?

OpenStudy (rulnick):

11^2 = x produces x=121. sqrt(225) = sqrt(15*15) = 15.

OpenStudy (unklerhaukus):

well the first is from your 11 times tables

OpenStudy (anonymous):

THANK You!!!!!!!!!!!!!!!!!!!